A 2kg block is released 4m from a massless spring with a force constant K=100N/m that is aong a plane inlined at 30°. if the plane is frictionless , find the maximum compression of the spring.

1
Expert's answer
2021-12-07T20:11:33-0500

"mgh\sin\alpha=\frac{kx^2}2,\implies"

"x=\sqrt{\frac{2mgh\sin\alpha}k}=0.89~m."